A loan that compounds interest monthly has an APR of 22 percent. What is the EAR?

a)24.36 percent

b)

23.89 percent

c)

22.00 percent

d)

23.12 percent
Answer #1

Formula for effective interest rate is:

r = (1+i/n) n – 1        

r = Effective interest rate (EAR)

i = Stated interest rate (APR) = 22 %

n = No. of compounding periods in a year = 12

r = (1+ 22 %/12) 12 – 1

= (1+ 0.22/12)12 – 1

= (1+ 0.0183333333333333)12 – 1

= (1.0183333333333333)12 – 1

= 1.24359657794443 – 1

= 0.24359657794443 or 24.36 %

EAR of the loan is 24.36 %

Hence option “a) 24.36 percent” is correct answer.